OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 10 = x2 + 3x? (−2, 0) and (−5, 0) (2, 0) and (5, 0) (−2, 0) and (5, 0) (2, 0) and (−5, 0)

OpenStudy (psymon):

have you learned much about quadratics and factoring yet?

OpenStudy (anonymous):

yes i need to make sure if the A is the correct

OpenStudy (anonymous):

sorry i mean D

OpenStudy (psymon):

Alrighty: y = x^2 + 3x - 10 (5)(-2) = -10 5 + -2 = 3 (x+5)(x-2) (2, 0) (-5,0) So yep, youre right, lol.